Find the area of each of the following.
1. square whose sides are (x+1) in length
2. triangle whose base is 5 1/2 in. and whose height is 2 1/4 in.
For #1, I tried it and got x^2+1
For #2, I got 6 3/16

Answer by Cromlix(4381)

Area of a square with sides (x + 1)
= (x + 1)(x + 1) = x^2 + 2x + 1
Area of a triangle whose base is 51/2ins
and height 21/4 ins.
= 1/2(51/2 + 21/4)
= 1/2 (31/4)
= 31/8
= 3 7/8 ins^2
Hope this heps.
:-)